OpenStudy (anonymous):

solve for x: 15+ (x+17 = 3x-10

OpenStudy (anonymous):

15+ x+17 = 3x-10 32+x=3x-10 32+x-x=3x-x-10 32=2x-10 32+10=2x 42=2x x=21

OpenStudy (anonymous):

really?? it's that easy??